The control room was filled with people tonight to witness the official first light tonight (as opposed to the dress rehearsal last Friday). So far everything has gone well other than the weather. Around 9:30 pm Krissy opened the dome (as she predicted after dinner), allowing us to take even more images of star clusters (mostly chosen to make focussing easy). Unfortunately, only half an hour later we had to close the dome again due to imminent thunderstorms and rain.
Nevertheless, we made progress on several fronts today:
- The ring we saw on Friday's image was most likely caused by condensation on the dewar window (Thanks for the suggestion, George J). This was resolved by increasing the dry air flow. After some time spent wondering, this has turned out to be no issue at all.
- The baffle to block a stray light path (from the tertiary mirror via the primary mirror into the instrument) was installed today. We tested the effect of the baffle, and on first look it seems to remove the predicted scattered light component.
- We saw a << 1% pupil ghost in r' band flat fields, but apparently in no other filter. We expect that Frank Valdes' & Rob Swater's pipeline can readily handle the ghost.
- Examining an 20 second exposure of a 2nd magnitude star, we have not seen any indication of strong crosstalk yet. However, before making a bold claim here, we will obtain more suitable observations to quantify any crosstalk.
- A first experiment indicates that we have no obvious light leaks in the instrument – detailed testing will follow soon.
Today marks the transition from instrument installation into the engineering commissioning phase. This week we will concentrate on fundamental detector operations, harvesting header information from the telescope, and enabling telescope guiding with ODI guide star videos. During this phase we might not produce as many spectacular images, but we will continue to keep you up to date on our progress.
Daniel & Todd
|More stars, nicely packaged into an open cluster (do you recognize it, Bob?). 30 seconds in r', bias and flat field corrected. Seeing was about 1"|